# Towards deep physics-informed Kolmogorov–Arnold networks

### v1

@@ Appendix E PirateNet Architecture
In this appendix, we provide a detailed description of the PirateNet architecture [23] employed in the benchmarks of Section 5, together with an explicit parameter count. As in Section 4.2, we consider a single input sample $\mathbf{x}\in\mathbb{R}^{d_{\text{I}}}$, where $d_{\text{I}}$ is the number of input coordinates. If periodic boundary conditions are present, they are first enforced through the embedding of Eq. (39) in the main text, resulting in the transformed input $\tilde{\mathbf{x}}\in\mathbb{R}^{\tilde{d}_{\text{I}}}$.
@@
All MLP layers, including the two of Eq. (107), follow the Random Weight Factorization (RWF) formulation [22], with weights initialized using the standard Glorot scheme [43] and biases initialized at zero.
@@
Spyros Rigas: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft. Fotios Anagnostopoulos: Data curation, Visualization, Writing – original draft. Michalis Papachristou: Data curation, Visualization, Writing – original draft. Georgios Alexandridis: Project administration, Supervision.

### v2

@@ Appendix G PirateNet Architecture
In this appendix, we provide a detailed description of the PirateNet architecture [PirateNets] employed in the benchmarks of Section 5, together with an explicit parameter count. As in Section 4.2, we consider a single input sample $\mathbf{x}\in\mathbb{R}^{d_{\text{I}}}$, where $d_{\text{I}}$ is the number of input coordinates. If periodic boundary conditions are present, they are first enforced through the embedding of Eq. (32) in the main text, resulting in the transformed input $\tilde{\mathbf{x}}\in\mathbb{R}^{\tilde{d}_{\text{I}}}$.
@@
All MLP layers, including the two of Eq. (113), follow the Random Weight Factorization (RWF) formulation [RWF], with weights initialized using the standard Glorot scheme [Glorot] and biases initialized at zero.
@@
Spyros Rigas: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft, Writing - Review & Editing. Fotios Anagnostopoulos: Data curation, Visualization, Writing – original draft. Michalis Papachristou: Data curation, Visualization, Writing – original draft. Georgios Alexandridis: Project administration, Supervision.